# Dynamic Benchmark Construction for Evaluating Large Language Models on Real-World Codes

### v1

@@ 6 Conclusion & Future Work
We introduced Code2Bench, a novel automated framework for dynamically constructing rigorous, contamination-resistant benchmarks from real-world code. By integrating dynamism, systematic dependency control, and Property-Based Testing, Code2Bench addresses limitations of existing benchmarks and enables reliable evaluation. Our benchmark, Code2Bench-2505, revealed that state-of-the-art LLMs struggle with complex logic (SC tasks) and cross-language generalization but perform better on common library interactions (WSC tasks in Python). These results highlight specific LLM weaknesses and demonstrate the value of Code2Bench in providing deeper insights into realistic coding challenges.
Despite these advances, our work has limitations in evaluation scope and task coverage, discussed in detail in Appendix J. Future work will focus on extending Code2Bench’s capabilities. Key directions include expanding language support by leveraging the framework’s language-agnostic core and enabling domain-specific benchmark construction. The language-agnostic Scope Graph-based dependency analysis , central to our framework, is inherently extensible. By defining Scope Graph rules for domain-specific libraries and APIs (e.g., system calls in systems programming or specialized data science libraries), Code2Bench can automatically generate tasks tailored to real-world coding challenges in these domains.
@@ A.1 Source Code Parsing via Abstract Syntax Trees (ASTs)
Following the initial identification of candidate functions from the source code repositories, the raw text code of these functions and their surrounding context is processed using Tree-sitter [29]. Tree-sitter is a parser generator tool that produces concrete syntax tree parsers for various programming languages. Unlike traditional compilers that focus on semantic analysis, Tree-sitter is specifically designed for source code analysis tools, providing robust, incremental parsing capabilities and generating detailed, well-structured Concrete Syntax Trees (CSTs), which are closely related to Abstract Syntax Trees (ASTs).
@@ C.2 Complexity Assessment via Cyclomatic Complexity
To focus the benchmark on tasks that require models to generate non-trivial code logic, we assess the structural complexity of each candidate function using Cyclomatic Complexity ($CC$) [31]. Cyclomatic Complexity is a quantitative measure of the number of linearly independent paths through a program’s source code. It is calculated based on the CFG of the function using the formula:
@@ Appendix D Semantic Filtering and Difficulty Assessment
This appendix provides additional technical details regarding the LLM-based Semantic Filtering process used in the Candidate Filtering stage (Section 2.1) to refine the candidate function set through semantic assessment. While the main text outlines the motivation and overall structure, here we present in detail the prompts employed during this stage of semantic filtering.
@@
- **Basic Types Only**: The function’s parameters must be basic types (e.g., ‘int‘, ‘float‘, ‘str‘, ‘list‘, ‘dict‘, etc.).
- If the parameters include custom classes or unknown types, the function is **not suitable**.
- If the function relies on methods or attributes of unknown objects, the function is **not suitable**.
@@
- **Meaningful Complexity**: The function should provide a meaningful test of the model’s capabilities.
- If the function is very long but its logic is trivial, it is **not suitable**.
- If the function is too simple (e.g., basic business logic), it is **not suitable**.
- If the function lacks clear logic or purpose, it is **not suitable**.
@@
- **No External Imports**: The function should not import other modules or depend on external libraries.
@@
“‘json
"Suitable": false,
"Reason": "The reason why the function is not suitable for generating an instruction."
@@
- **General Applicability**: The function should not require highly specialized domain knowledge to understand or implement.
- If the function is tied to a specific, non-generalizable application (e.g., proprietary system logic), it is **not suitable**.
@@ Appendix E Property-Based Testing
This appendix provides additional technical details regarding the Property-Based Testing (PBT) process used in the Benchmark Instance Construction stage (Section 2.2 ) to generate rigorous test cases for Code2Bench. While it describes the core principles, here we elaborate on the technical implementation.
@@ E.4 Code Example (Python/Hypothesis)
import hypothesis.strategies as st
from hypothesis import given, settings, Verbosity
[…]
# Strategy for generating arbitrary strings
string_strategy = st.text()
[…]
return s[::-1]
[…]
# Hypothetical LLM output, might have bugs
return "".join(reversed(s))
[…]
@given(s=string_strategy) # Use the defined strategy to generate input s
@settings(max_examples=1000) # Example settings for testing
def test_reverse_property(s):
# Property: Reversing twice returns the original string
assert $reverse_string($s)) == s # This property tests the function against itself
[…]
expected_output = ground_truth_reverse(s)
actual_output = llm_generated_reverse(s)
assert actual_output == expected_output # This property tests LLM output against GT
[…]
@@
E.5 Ensuring Test Case Validity and Cross-language Portability
This section elaborates on how Code2Bench ensures the generated test cases are both valid for evaluation and portable across different target languages for SC tasks, crucial aspects of our rigorous testing methodology.
Ensuring Validity via Ground Truth Validation
A key step in test case generation is ensuring that the inputs generated by PBT strategies are valid and do not cause unexpected errors in the ground truth implementation. While strategies aim to generate diverse inputs, including edge cases, some inputs might be fundamentally invalid according to the function’s inherent constraints or assumptions (e.g., negative values for a parameter that must be non-negative, or strings in an unexpected format). We execute the ground truth function with each generated input. If the ground truth throws an expected exception (e.g., “ValueError” for invalid input according to the function’s contract) or any unexpected runtime error, that input is discarded. Only inputs that the ground truth can process without crashing, producing a valid output, are included in the final test suite. This process, illustrated in the code example below, guarantees that all test cases in the benchmark are meaningful and that failures of the LLM’s generated code are due to functional incorrectness rather than invalid inputs that crash the ground truth.
Ensuring Cross-language Portability for SC Tasks
For Self-Contained (SC) tasks, the generated test suites are designed to be reusable for evaluating LLMs across multiple programming languages (Python, Java, Go, TS, JS). To achieve this, the Strategy Builder component enforces constraints on the generated inputs to ensure they use basic data types (integers, floats, strings, lists, dictionaries, booleans) with ranges safely representable across these common target languages. For instance, instead of allowing Hypothesis to generate Python’s arbitrary precision integers, explicit “min_value” and “max_value” bounds are set for integer strategies to fit within standard 32-bit or 64-bit integer limits common in Java, Go, and TypeScript. Similarly, constraints are applied to float ranges and the complexity of nested structures. This focus on portable basic types and bounded ranges ensures that the test case data can be correctly loaded and processed by test runners in different languages without issues like integer overflow or incompatible data structures, making the evaluation truly cross-lingual for SC tasks.
Code Example
Below is a simplified code example illustrating the core PBT generation logic, including calling the ground truth and handling exceptions to ensure test case validity.
import hypothesis.strategies as st
from hypothesis import settings, given
# import other necessary libraries
[…]
# Might raise exceptions for invalid inputs
def example_ground_truth(input1, input2):
# implementation logic
pass
[…]
def strategy_for_input1():
# strategy definition for input1
pass
[…]
# strategy definition for input2
# Ensure ranges are portable for SC tasks!
pass
[…]
# @settings(…)
@given(
input1=strategy_for_input1(),
input2=strategy_for_input2()
def test_generate_valid_cases(input1, input2):
# 1. Validate input and Compute expected output using the ground truth
try:
# Call ground truth with generated inputs
expected_output = example_ground_truth(input1, input2)
except Exception as e: # Catch any exception from ground truth
# Discard inputs that caused errors in the ground truth
return # Hypothesis discard mechanism
[…]
# This pair is collected and saved for the benchmark.
test_case = {"Inputs": {"input1": input1, "input2": input2}, "Expected": expected_output}
collection_list.append(test_case)
pass # Placeholder for collection logic
[…]
This example code snippet demonstrates the core mechanism of validating generated inputs by running the ground truth and collecting only those inputs that result in a valid output, thereby ensuring test case validity.
@@ Appendix F Testcase Runner Generation
This appendix provides a detailed presentation of the prompts used for generating test runners (Section 2). Since test runners are inherently language-specific, we designed tailored prompts for each programming language. The corresponding prompts are listed in Tables 10 through 13.
Table 10: Prompt Template for TS Testcase Runner Generation in Code2Bench
# System
## Task Description
As an expert TypeScript developer specializing in test case generation and function signature translation, your task is to generate a function signature based on a Python function and its test cases.
### Requirements
1. **Function Signature**:
- Match the Python function’s input and output types.
- Use export function declarations, include a // TODO: Implement comment inside the function.
- Output in ‘<signature>‘ tags.
2. **Special Considerations**:
- Use explicit and specific TypeScript types, avoiding any or object for type safety.
- Handle complex inputs (e.g., arrays, tuples, nested structures) precisely.
- Return only ‘<signature>‘ sections, without additional explanations.
## Input Format
- **Python Function**: A Python function to be tested, including its signature and implementation, provided as code.
## Output Format
“‘plaintext
<signature>
[TypeScript function signature]
</signature>
## Examples
### Example 1: fast_format_html
#### Input
Python Function:
“‘python
def fast_format_html(html_string):
indent = 0
indent_str = " "
formatted = []
[…]
for part in parts:
if not part.strip():
continue
if part.startswith("</"):
indent -= 1
formatted.append(indent_str * indent + part)
elif part.startswith("<") and part.endswith("/>"):
formatted.append(indent_str * indent + part)
# User
The previously generated runner code:
[Runner Code]
[Error Message]
[Function Message]
Table 11: Prompt Template for Java Testcase Runner Generation in Code2Bench
@@
Table 12: Prompt Template for Go Testcase Runner Generation in Code2Bench
# System
## Task Description
As an expert Go developer, generate two files from Python function definitions and JSON test cases:
1. ‘runner_test.go‘ - Test execution logic only
2. ‘tested.go‘ - Implementation signature with all required types
### Strict Requirements
1. **File Separation Rules**:
2. **Error Handling Rules**:
- If Python function can raise exceptions:
→ Go function MUST return ‘(resultType, error)‘
→ Test MUST verify no error occurs (since test cases are all valid inputs)
- If Python has no error cases:
→ Go function returns only ‘resultType‘
3. **Type Definition Rules**:
| Usage Scenario | Location | Example |
| Used in function signature | ‘tested.go‘ | ‘type TagInfo struct{}‘ |
| Only in test JSON structure | Inline in TestCase | ‘Inputs struct{ Temp float64 }‘ |
4. **Test Validation Logic**:
- For functions returning ‘(result, error)‘:
### Output Format
“‘plaintext
<code>
// runner_test.go
[Complete test file content]
</code>
<signature>
// tested.go
[Complete implementation file content]
</signature>
## Examples
## Note
- Ensure the generated runner_test.go file is complete and executable.
- The function signature should be a valid Go function signature that matches the Python function’s behavior.
- Only return the runner_test.go file in ‘<code>‘ and the tested.go in ‘<signature>‘ tags.
# User
The previously generated runner code:
[Runner Code]
[Error Message]
[Function Message]
Table 13: Prompt Template for Weakly Self Contained Testcase Runner Generation in Code2Bench
@@ Appendix G Instruction Generation
Table 14: Prompt Template for Weakly Self Contained Instruction Generation in Code2Bench
@@
You are a python programming expert who is refining docstrings in existing programs. You will be given a python function in a python file with an existing (possibly underspecified) docstring with corresponding some input-output examples extracted. Your goal is to refine the associated docstring by making it more informative, precise and complete without adding verbosity or detailed programming logic to the docstring. When there is a docstring, the docstring is used to evaluate the code generation capabilities of a model.
@@
### Output Format - Return the docstring in <docstring> tags following the Google-style format.
@@
Table 15: Prompt Template for Self Contained Instruction Generation in Code2Bench
@@
- Do **not** add or modify examples from the ‘Example Usages‘ data.
@@ H.1 Benchmark Details
This subsection provides additional details regarding the construction and properties of the Code2Bench-2505 instance.
Source Code Acquisition and Filtering
The real-world code used for constructing Code2Bench-2505 was sampled from GitHub repositories with at least 500 stars. This criterion was applied to focus on relatively mature and well-maintained projects, increasing the likelihood of finding high-quality and realistic code functions. A comprehensive list of the specific GitHub repositories sampled for Code2Bench-2505 is provided alongside the benchmark data and code release at the public URL (as mentioned in abstract).
LLM Usage in Benchmark Construction
Large Language Models play a role in several stages of the Code2Bench pipeline for constructing benchmark instances. For the construction of Code2Bench-2505, GPT-4o was primarily employed for these tasks within the benchmark construction pipeline. It is crucial to emphasize that GPT-4o was strictly excluded from the set of models evaluated in the Evaluation section.
Quality Assurance and Release Process
Prior to the official release of any Code2Bench benchmark instance, a rigorous quality assurance process is performed. This includes a dry run using a designated, capable LLM. Given that the testcase runners for Python, JavaScript (JS), and TypeScript (TS) are primarily constructed using template-based methods, their inherent correctness is highly reliable. Consequently, the dry run specifically emphasizes validating the testcase runners for Java and Go, as well as the WSC Python runner
H.2 Benchmark Task Examples

### v2

@@ 6 CONCLUSION & FUTURE WORK
We introduced Dual Scaling, a new philosophy for benchmark construction, and presented CODE2BENCH, a framework that operationalizes it by systematically scaling the source of problems from real-world code and the rigor of tests via high-coverage PBT. Our evaluation on the resulting CODE2BENCH-2509 benchmark provided a deep, diagnostic analysis of modern code LLMs, revealing a consistent gap between their API application (WSC) and algorithmic synthesis (SC) capabilities, and quantifying for the first time how language paradigms shape their failure modes.
Limitations and Future Work.
Our work, while establishing a robust framework for evaluating functional correctness, opens several avenues for future expansion. We plan to extend our Scaling the Source principle to repository-level, Project-Dependent (PD) tasks to assess codebase understanding. Concurrently, we will expand our Scaling the Rigor principle to incorporate non-functional properties such as code efficiency and security. By evolving along these axes, CODE2BENCH will continue to provide a challenging and realistic measure of true software engineering competence.
Reproducibility Statement
We are committed to ensuring the full reproducibility of our work. The complete CODE2BENCH-2509 benchmark suite, including all task instructions, ground-truth solutions, and PBT-generated test suites scripts, is also included. The project, including all data and results, is also available at our anonymized repository: https://code2bench.github.io/.
@@ A.1 Source Code Parsing via Abstract Syntax Trees (ASTs)
Following the initial identification of candidate functions from the source code repositories, the raw text code of these functions and their surrounding context is processed using Tree-sitter tree-sitter (2024). Tree-sitter is a parser generator tool that produces concrete syntax tree parsers for various programming languages. Unlike traditional compilers that focus on semantic analysis, Tree-sitter is specifically designed for source code analysis tools, providing robust, incremental parsing capabilities and generating detailed, well-structured Concrete Syntax Trees (CSTs), which are closely related to Abstract Syntax Trees (ASTs).
@@ C.2 Complexity Assessment via Cyclomatic Complexity
To focus the benchmark on tasks that require models to generate non-trivial code logic, we assess the structural complexity of each candidate function using Cyclomatic Complexity ($CC$) 2. Cyclomatic Complexity is a quantitative measure of the number of linearly independent paths through a program’s source code. It is calculated based on the CFG of the function using the formula:
@@ Appendix D Semantic Filtering and Difficulty Assessment
This appendix provides additional technical details regarding the LLM-based Semantic Filtering process used in the Candidate Filtering stage to refine the candidate function set through semantic assessment. While the main text outlines the motivation and overall structure, here we present in detail the prompts employed during this stage of semantic filtering.
To validate the robustness and objectivity of our LLM-based filtering, we conducted two inter-rater reliability studies. Inter-LLM Agreement. We measured the agreement between three diverse, state-of-the-art LLM judges (GPT-4o, Claude-4-Sonnet, and a Qwen3-Max) on a random sample of 100 function candidates. For the binary task of semantic filtering (trivial vs. meaningful), the judges achieved a Fleiss’ Kappa of 0.92, indicating almost perfect agreement. Human-LLM Agreement. We further compared the judgments of our primary LLM judge (GPT-4o) against a consensus gold standard established by two human experts with extensive programming experience. On a set of 50 functions, the analysis yielded a Cohen’s Kappa of 0.95.
@@
- **Basic Types Only**: The function’s parameters must be basic types (e.g., ‘int‘, ‘float‘, ‘str‘, ‘list‘, ‘dict‘, etc.)…
@@
- **Meaningful Complexity**: The function should provide a meaningful test of the model’s capabilities…
@@
- **No External Imports**: The function should not import other modules or depend on external libraries…
@@
- **General Applicability**: The function should not require highly specialized domain knowledge to understand or implement…
@@ Appendix E Property-Based Testing
This appendix provides additional technical details regarding the Property-Based Testing (PBT) process to generate rigorous test cases for Code2Bench. While it describes the core principles, here we elaborate on the technical implementation.
@@ E.4 Code Example (Python/Hypothesis)
1import hypothesis.strategies as st
2from hypothesis import given, settings, Verbosity
3
4# 1. Define strategies for input generation
5# Strategy for generating arbitrary strings
6string_strategy = st.text()
7
8def ground_truth_reverse(s: str) -> str:
9 return s[::-1]
10
11def llm_generated_reverse(s: str) -> str:
12 # Hypothetical LLM output, might have bugs
13 return "".join(reversed(s))
14
15# Define a property using @given decorator
16@given(s=string_strategy) # Use the defined strategy to generate input 's'
17@settings(max_examples=1000) # Example settings for testing
18def test_reverse_property(s):
19 # Property: Reversing twice returns the original string
20 assert $reverse_string($s)) == s # This property tests the function against itself
21
22 # Property: LLM output matches Ground Truth output
23 expected_output = ground_truth_reverse(s)
24 actual_output = llm_generated_reverse(s)
25 assert actual_output == expected_output # This property tests LLM output against GT
@@ Appendix F Testcase Runner Generation
This appendix provides a detailed presentation of the prompts used for generating test runners (Section 2). Since test runners are inherently language-specific, we designed tailored prompts for each programming language. The corresponding prompts are listed in Tables 9 through 10.
Table 9: Prompt Template for SC-Java Testcase Runner Generation in Code2Bench
@@
Table 10: Prompt Template for WSC-Python Testcase Runner Generation in Code2Bench
@@ Appendix G Instruction Generation
Table 11: Prompt Template for WSC-Python Instruction Generation in Code2Bench
@@
You are a python programming expert who is refining docstrings in existing programs. You will be given a python function in a python file with an existing (possibly underspecified) docstring with corresponding some input-output examples extracted.
Your goal is to refine the associated docstring by making it more informative, precise and complete without adding verbosity or detailed programming logic to the docstring. When there is a docstring, the docstring is used to evaluate the code generation capabilities of a model.
@@
### Output Format
- Return the docstring in <docstring> tags following the Google-style format.
@@
Table 12: Prompt Template for SC-Python Instruction Generation in Code2Bench
@@
- Preserve all original examples **verbatim** in the final docstring’s ‘Examples‘ section…
@@
Table 13: Prompt Template for SC-Java Instruction Generation in Code2Bench
# System
You are an expert Java architect and technical writer, specializing in creating high-quality, professional Javadoc documentation. Your task is to generate a precise and informative Javadoc for a given Java method, enabling another senior Java developer to re-implement it accurately within a complete, self-contained ‘Tested.java‘ file.
## Core Objective
The generated Javadoc and method signature must serve as a perfect "specification" for the provided ground-truth method.
## Requirements
1. Core Description Fidelity: The Javadoc must accurately reflect the method’s behavior, including the precise sequence of checks, conditions, and resulting actions within the code.
2. Edge Cases: Detail how the method handles edge cases, such as null, empty, or special/magic values.
3. Data Structures: Accurately describe parameters and return values using standard Java types.
4. Javadoc Refinement: If the method already has a Javadoc, your primary goal is to refine and enhance it to meet these high standards. Integrate existing accurate information with your new insights. Do not discard valuable details from the original author.
5. Example Handling:
If the original Javadoc contains an example (e.g., in a ‘<pre>@code …</pre>‘ block): Preserve the original example verbatim.
If not: Omit any example section entirely.
## **Input Structure**
“‘java
{ground_truth_java_method_code}
## Output Structure
Return a single ‘<signature>‘ tag containing a complete, self-contained, and runnable ‘Tested.java‘ file content. The content must be enclosed in a “‘java‘ code block and include all necessary imports, the generated Javadoc, the ‘public class Tested‘, and the public static method signature with a ‘// TODO: implement this method‘ comment.
<signature>
“‘java
// All necessary imports (e.g., java.util.*) should be here.
import java.util.List;
import java.util.Map;
public class Tested {
High-quality, Google-style Javadoc goes here. …
public static ReturnType methodName(ParameterType parameterName) {
// TODO: implement this method
</signature>
## Final Instructions
- Ensure the method signature in the ‘<signature>‘ tag perfectly matches the ground truth, including visibility (‘public static‘), return type, method name, and parameter types.
- Don’t add implementation details.
- Include all necessary imports based on the types used in the signature.
# User
<function>
[Function Code]
</function>
@@ H.1 Benchmark Details
This appendix provides a detailed breakdown of the construction process and diversity analysis for the CODE2BENCH-2509 suite. All data was sourced from public GitHub repositories with commits made between May 2025 and September 2025.
To ensure the diversity, quality, and representativeness of the source code used in Code2Bench, we adhered to a strict, multi-stage data selection protocol. This appendix details the criteria and sampling strategies employed during the “Scaling the Source” phase.
We targeted high-quality, actively maintained, and well-tested open-source repositories hosted on GitHub. To be included in our initial candidate pool, a repository must meet the following quantitative thresholds:
  • Community Validation: $\geq 500$ Stars. This threshold filters out personal experiments and ensures a baseline of community scrutiny and adoption.
  • Active Maintenance: At least one commit within the 3 months prior to our data collection cutoff (May 2025). This ensures the code reflects modern coding practices and library versions.
  • Test Availability: Must contain an identifiable test suite (e.g., presence of tests/ directory, usage of pytest/junit). This is crucial for verifying our ground truth extraction.
  • License Permissibility: Must be under permissive licenses (e.g., MIT, Apache 2.0, BSD) to allow for redistribution and modification.
To mitigate noise and bias, we applied semantic filters to exclude repositories that do not represent real-world software engineering contexts:
  • Homework & Tutorials: Repositories with keywords like “assignment”, “course”, “tutorial”, “learn-python/java”, or “leetcode” in their description or README were excluded. These typically contain toy problems distinct from production-grade code.
  • Aggregators & Forks: We excluded repositories identified as mere collections of other projects (e.g., “awesome-xxx”) or direct forks without significant divergence, ensuring unique data sources.
To prevent domain bias (e.g., over-representation of web frameworks), we employed a stratified sampling strategy based on GitHub topics and repository tags. We categorized candidates into 10 primary domains covering the spectrum of software development:
  1. 1.
    Web Development (e.g., frameworks, clients)
  2. 2.
    Data Science & Machine Learning (e.g., analytics, pipelines)
  3. 3.
    System Utilities (e.g., cli tools, file manipulation)
  4. 4.
    Network & Protocol (e.g., async io, sockets)
  5. 5.
    Security & Cryptography
  6. 6.
    Database & Storage
  7. 7.
    Text Processing & NLP
  8. 8.
    Media & Graphics (e.g., image processing)
  9. 9.
    Scientific Computing
  10. 10.
    Development Tools (e.g., linters, parsers)
We performed random sampling within each stratum to build our final source list of Python and Java repositories.
To fully support reproducibility and auditability, the complete list of source repositories is available on our open-source project page: https://code2bench.github.io/.
H.2 The Data Funnel: From Raw Functions to a Gold Standard
The final size of our benchmark is a direct result of a stringent, multi-stage filtering pipeline designed to prioritize quality, realism, and rigor. Table 14 illustrates this "Great Filter" process for the Python components, starting from over one million recently updated functions identified in 220 repositories. This rigorous process ensures that only the most suitable and high-quality candidates become part of the final benchmark.
Table 14: The Data Funnel for CODE2BENCH-2509 Python components, illustrating the multi-stage filtering process that prioritizes quality and rigor.
Stage Filtering Action & Criteria SC Candidates WSC Candidates
1. Initial Pool Functions parsed from recent commits ~1.17 Million
2. Dependency Filter Scope Graph: Strictly SC / WSC compliant 27,649 12,335
3. Testability/Complexity Testable outputs & Cyclomatic Complexity [2,10] 7,102 3,278
4. Sub-sampling Breadth-first sampling for diversity 901 432
5. Semantic Filter LLM-as-a-judge removes trivial tasks 479 315
6. The Great Filter PBT: 100% Branch Coverage Guarantee 217 194
H.3 Task Diversity Analysis
The high quality of CODE2BENCH-2509 is rooted in its rich diversity of tasks and application domains.
H.3.1 SC-Python: Algorithmic and Real-World Logic
The 217 tasks in the SC-Python component cover a wide spectrum of real-world programming challenges beyond simple puzzles. Key functional categories include:
  • Text Processing & Formatting: Ranging from LaTeX sanitization (strip_latex) to complex wrapping for SVG elements (wrap_text_for_svg).
  • Classic & Modern Algorithms: Includes fundamental algorithms like levenshtein and binary_search, as well as logic relevant to modern development tools like parsing LCOV reports (parse_lcov).
  • Parsing & Extraction: Challenges models to parse structured data from unstructured text, such as extracting JSON from noisy strings (_extract_balanced_json) or parsing version numbers (parse_version).
  • AI/LLM-Specific Logic: A unique feature is the inclusion of tasks from the AI ecosystem, such as formatting LLM error messages (format_llm_error_message) and parsing model outputs (extract_weave_refs_from_value).
  • Complex Data Structure Manipulation: Requires deep understanding of nested structures (e.g., flatten_state_dict, deep_merge).
H.3.2 WSC-Python: A Broad and Realistic API Ecosystem
The 194 tasks in the WSC-Python component require the use of over 35 distinct libraries and modules [^1] This count is based on unique top-level import statements, e.g., re, numpy, scipy.spatial., ensuring a faithful evaluation of a model’s practical API fluency. The distribution is representative of real-world Python development, covering a wide and diverse ecosystem rather than focusing on a narrow set of APIs. The required libraries span multiple key domains of software engineering:
  • Data Processing & Text Manipulation: A significant portion of tasks involve core data handling using standard libraries like re, json, ast, datetime, urllib, unicodedata, and base64.
  • Scientific Computing & Data Science: The benchmark probes capabilities in specialized numerical and data-centric domains, requiring libraries such as numpy, scipy (including submodules like scipy.spatial.transform), and pandas.
  • Machine Learning: Uniquely, our suite includes tasks that interact with the machine learning ecosystem, leveraging modules from scikit-learn like TfidfVectorizer and cosine_similarity.
  • Advanced Standard Library Proficiency: Beyond common utilities, the tasks require a deep knowledge of Python’s standard library, including advanced modules such as itertools, collections (e.g., Counter, defaultdict), difflib, bisect, and struct.
This broad and realistic library coverage ensures that WSC-Python provides a holistic assessment of a model’s ability to function as a practical coding assistant in a diverse range of real-world scenarios.
H.3.3 SC-Java: Validating Extensibility with Diverse Tasks
The successful generation of the 249-task SC-Java suite provides concrete evidence of our framework’s extensibility. The quality of this component is validated by its high diversity, which mirrors the real-world complexity found in its Python counterpart. The tasks span a wide array of application domains:
  • String Manipulation & Parsing: A large portion of tasks involve complex string operations, such as format conversion (convertToCamelCase), cleaning (cleanText), and escaping for different contexts (escapeJsonString, escapeCsvField).
  • Encoding & Data Conversion: Numerous tasks focus on byte-level manipulation, primarily converting between byte arrays and hexadecimal strings (e.g., bytesToHex), a common task in systems programming and networking.
  • Mathematics & Algorithms: The suite includes non-trivial algorithms like checksum validation (luhnBankCardVerify) and edit distance calculation (levenshteinDistance).
  • Domain-Specific Logic: Crucially, the tasks are not generic puzzles but are rooted in specific application domains, including game development (e.g., Minecraft metadata transformation, transformMetaDecoModel) and systems utilities (e.g., calculating CPU affinity masks, maskToCpuAffinity).
This demonstrates our framework’s ability to extract meaningful and realistic algorithmic challenges from any complex, real-world codebase, regardless of the programming language.
H.4 Benchmark Task Examples